Dakota Company experienced the following events during 2018:

  1. Acquired $15,000 cash from the issue of common stock.

  2. Paid $15,000 cash to purchase land.

  3. Borrowed $5,000 cash.

  4. Provided services for $55,000 cash.

  5. Paid $2,000 cash for utilities expense.

  6. Paid $45,000 cash for other operating expenses.

  7. Paid a $1,000 cash dividend to the stockholders.

  8. Determined that the market value of the land purchased in Event 2 is now $16,000.

Required

  1. a. The January 1, 2018, general ledger account balances are shown in the following accounting equation. Record the eight events in the appropriate general ledger accounts. Record the amounts of revenue, expense, and dividends in the Retained Earnings column. Provide the appropriate titles for these accounts in the last column of the table. The first event is shown as an example.

  2. b-1. Prepare an income statement for the 2018 accounting period.

  3. b-2. Prepare a statement of changes in equity for the 2018 accounting period.

  4. b-3. Prepare a year-end balance sheet for the 2018 accounting period.

  5. b-4. Prepare a statement of cash flows for the 2018 accounting period.

  6. c. Determine the percentage of assets that were provided by retained earnings. Can you determine the cash in retained earnings?

Dakota Company experienced the following events during 2018: 1. Acquired $15,000 cash from the issue of common stock. 2. PaidComplete this question by entering your answers in the tabs below. Req A Req B1 Req B2 Req B3 Req B4 Reqc The January 1, 2018Req B1 Complete this question by entering your answers in the tabs below. Rega Req A Req B1 Req B1 Req B2 Req B3 Req B4 ReqcComplete this question by entering your answers in the tabs below. Req A Reg A Req B1 Req B1 Req B2 Req B3 Req B4 Reqc PreparComplete this question by entering your answers in the tabs below. Req A Req B1 Req B2 Req B3 Req B4 Reqc Prepare a year-endPrepare a statement of cash flows for the 2018 accounting period. (Amounts to be deducted should be indicated with a minus siComplete this question by entering your answers in the tabs below. Req A Req B1 Req B2 Req B3 Req B4 Reqc Determine the perce
Answer #1

A)

DAKOTA COMPANY
Accounting Equation for 2018
Event Assets = Liabilities + Stockholders' Equity Account titles for Retained Earnings
Cash + Land = Notes Payable + Common Stock + Retained Earnings
Bal. 1/1/2018 $14500 + 10000 = 10000 + 10000 + 4500
1. 15000 + = + 15000 + NA
2. -15000 + 15000 = + + NA
3. 5000 + = 5000 + + NA
4. 55000 + = + + 55000 Service Revenue
5. -2000 + = + + -2000 Utilities expense
6. -45000 + = + + -45000 Operating expense
7. -1000 + = + + -1000 Dividends
8. + = + + NA
Totals $26500 + 25000 = 15000 + 25000 + 11500

The market value of the land does not have any effect in the accounting equation.

B-1)

DAKOTA COMPANY
Income Statement
For the Year Ended December 31, 2018
Service Revenue $55000
Utilities expense -2000
Operating expense -45000
Net income $8000

B-2)

DAKOTA COMPANY
Statement of Changes in Stockholders' Equity
For the Year Ended December 31, 2018
Beginning common stock $10000
Plus: Common stock issued 15000
Ending common stock $25000
Beginning retained earnings $4500
Plus: Net income 8000
Less: Dividends -1000
Ending retained earnings 11500
Total stockholders' equity $36500

B-3)

DAKOTA COMPANY
Balance Sheet
As of December 31, 2018
Assets
Cash $26500
Land 25000
Total assets $51500
Liabilities
Notes payable 15000
Total liabilities 15000
Stockholders' Equity
Common stock 25000
Retained earnings 11500
Total stockholders' equity 36500
Total liabilities and stockholders' equity $51500

B-4)

DAKOTA COMPANY
Statement of Cash Flows
For the Year Ended December 31, 2018
Cash flows from operating activities:
Received from customers $55000
Utilities expenses paid -2000
Operating expenses paid -45000
Net cash flow from operating activities $8000
Cash flows from investing activities:
Purchase of land -15000
Net cash flow from investing activities -15000
Cash flows from financing activities
Issuance of common stock 15000
Cash borrowed 5000
Dividend paid -1000
Net cash flow from financing activities 19000
Net increase in cash (8000-15000+19000) 12000
Add: Beginning cash balance 14500
Ending cash balance $26500

C) Percentage of assets= Retained earnings*100/Assets

= $11500*100/51500= 22.3%

Can you determine the cash in retained earnings?

No, cash cannot be determined from the retained earnings.